Telescope/TCS Checkout for LUCIFER Com3
UT 2009 January 09

Observers: D.Thompson (@LBT) with N.Ageorges, W.Seifert, R.Gredel, M.Juette plus A. Rakich and J. Hill in Tucson for part of night.
Telescope Operator: A.Ceranski
SW Support: P.Grenz in Tucson

Summary:

  • ADD 3.75mm to the nominal AGW focus in the config file.
  • done the following afternoon by Torsten. now 35.35, was 31.6.

GCS WFSC /Repository/AGW_Data/090109w_gcs and field data in 090109w (DLM)

Guider images in /tmp/20090109 (DLM)

Details:

17:18 Not photometric. There are thick cirrus around, should clear off later. T=5.6, RH=15.5, Wind ~6m/s average.

17:45 Dome open

17:53 Aaron has finished encoder resets. Still having problems with the fiducial of encoder 0, but telescope coordinates should be OK.

Startup pointing & collimation correction

pointto M5_1040 gs=0 ACQUIRE PAM- (don't see the star, but clouds are fairly heavy)
18:00 pointto M5_1568 gs=0 ACQUIRE PAM-
18:05 pointto M5_1568 TRACK
18:06 pointto M5_1568 TRACK PAM-
IE changed from 98 to 24.3 and CA from 75 to 53.75 (~90 arcsec offset)

Preset to BS9105 (EL~57), not found (looking through some thick clouds)
18:20:03 Preset to ACT0227 nearby to check. Preset hung. Killed GCS. This frees LUCIFER from the hang. GCS was involved because it had to park the probe. It seems to have done so, but did not communicate back to IIF that it was done?

Thu Jan 8 18:26:31.038 2009 setting none mode on left side
RPC error on MCSPU, additional problems.

Collimating on BS9105 while the clouds break up Thu Jan 8 19:05:45.548 2009 setting track mode on left side

07:11 Repeat 3a and 3b tests from last night

To hold WFS corrections, on Left PSF Wavefront Control GUI, do:
ButtonSorted ascending 3a 3b
Hold WFS Corrections yes no
Sec Act Opt Offload yes yes

3a Set the Z11 and Z22 gains in PSF to 0. Allow system to run for a while and then send a stepfocus. Note the image quality after the stepfocus. This will eliminate the small focus corrections that trigger the updates. But you probably don't want to run like this for more than an hour or so without a Z11 correction. The gains are changed by sending a manual correction of Z11 with the PSF Primary GUI. The manual corrections also allow changes of GAIN and FCIN along with the Zernike amplitude.

X=-1.601 Y=-0.378 Z=0.979 RX=-32.858 RY=28.690 EL=63:08 at 7:13
07:17 M1 total collimation still the same. WFS sees -264nm on Z4
07:18 sent stepfocus 0.0 Guider image looks good, but there was for 1 cycle an offset
07:19 First WFS image after stepfocus shows +246 on Z4
07:24 gains reset to 1
07:29 Z4 built up to -506 (670rms), sending stepfocus. FWHM on guider went from 0.7 to 0.5
07:31 Z4 = +433 on first WFS image after the stepfocus
07:33 Hold WFS corrections -> NO, allow to collimate

3b Set the FCIN flags Z11 and Z22 to 1. This causes the focus correction to be implemented by bending the mirror rather than by repositioning. In stable conditions using it this way will be no problem, and you will get continuous Z11 and Z22 corrections. However, in conditions with rapidly changing mirror temperatures there is a problem that the mirror will run out of its allowed range of correction forces and panic. Bending in the focus uses more force than getting focus by moving the mirror.

07:36 FCIN flags set to 1 for Z11 and Z22
X=-1.707 Y=-0.250 Z=0.845 RX=-32.047 RY=28.833
X=-1.710 Y=-0.245 Z=0.826 RX=-32.048 RY=28.746

Why did the primary move? Because we had not held the Z11/Z22 updates correctly.

07:43 sending stepfocus 0.0
07:45 Sety FCINs to 1 (last time they did not take).
(guide correction data from this test can be seen as an attachment to Issue #1904.)
X=-1.601 Y=-0.293 Z=0.835 RX=-32.609 RY=27.447
07:54 M1 did not move, corrections remained low. guide star seemed much more stable.

08:08 Preset to a PR/science target (NGC 1569)
had to correct pointing (IE=-5.8, CA=32.5), collimating off axis.

08:15 stepfocus 0.0 and Z11/Z22 gains to 0. SecActOptOffload = yes and Hold WFS Corrections = no
Taking a dithered set of data with LUCIFER.
X=-1.129 Y=-0.116 Z=1.075 RX=-29.611 RY=23.629

luci_20090108_0001-0100 (Z11/Z22 not applied at all)

08:42 Taking sky frames for the data above.

luci_20090108_0101-0131 Aborted script because an ABS offset was sent by mistake.

resent preset

luci_20090108_0132-0231 Sky images, effectively track mode. Stepfocus sent after each dither but not during the 20x5sec integrations.

Gains on Z11/Z22 reset to defaults, re-collimating

09:21 After recollimating, sent stepfocus to offload and changes FCINs to 1

luci_20090108_0232-0331 (Z11/Z22 applied as bending forces)

Significant jump visible in LUCIFER image 294, UT=04:36:27.6 midpoint of the 5 second image.

It also looks like the guider takes its first image after an offset a little early. Check timing issues.

9:51 Preset to astrometric field
Acquisition failed (image 000085). Source 18K above background but at edge. Check why this failed.
IE changed from -5.8 to -9.4, CA from 32.5 to 39.7
09:55 Resent preset...succeeded (collimation way out)

Measure/Correct LUCIFER rotation angle (15-30 minutes/field):

8 images with successful WCS. Offset is 0.243 degrees (rms = 0.045 degrees). LEFTZEROPOINT changed in PCSInstrument.conf to -187.245. Should recheck on multiple fields in coming nights.

BS9122 - Field Aberration Measurements data currently from last night
GStar X Y PA Doug's Angle wfscimage000*
0 3.463 612.288 0.00 -194 2-16
6 2.178 414.993 295.03 -499 17-19
0 3.463 612.288 295.03 -502 21-22
24 2.403 466.232 349.12 -199 23-25
0 3.405 612.335 349.12 -200 26-28
Data above in AGW_Data/090108w. Data below in AGW_Data/090109w
0 3.346 612.286 0 -506 2-8
6 2.168 414.916 295 -444 9-11
0 3.185 612.398 295 -444 12-14
24 2.324 466.219 349.12 -502 15-17
0 3.405 612.322 349.12 -504 18-20
46 2.137 424.475 93.0 -251 21-23
0 3.256 611.988 93 -254 24-27
didnt find star, going back on axis with PA 93
0 3.315 611.988 93 -259 28-31
29 2.718 442.987 179.79 -347 32-34
0 3.037 612.093 179.79 -350 35- 36
14 2.454 455.01 225.62 -398 37-39
see log entry below for 00:03
0 3.009 612.295 225.62 -414 40- 41
12 2.507 473.188 265.01 -456 42-44
0 3.053 612.429 265.01 -459 45- 47
6 2.208 414.981 295.03 -491 48-50
0 3.244 612.461 295.03 -493 51-52
24 2.391 466.231 349.12 -190 53-55
0 3.42 612.334 349.12 -191 56-59
46 2.137 424.406 93.0 -299 60-62
0 3.241 611.919 93.0 -300 63- 64
42 2.308 465.149 147.49 -3571 65-67
0 3.036 611.965 147.49 -359 68- 70

00:03 GCS failed guiding twice in the same way; acquired the guide star, overcorrected in the second guide image, then "star verification failed". Killed and restarted GCS and it happened again. see GCS log at time ~ 00:01:49 for the second instance.

Re-Measure AGW-PCS Transformation (2hr per field): see attached file AGWcoords.txt for data)

*BS9122 - AGw-PCS coordinate transformation calibration*
All data taken at PA=295.00
See attached file AGWcoords.txt for the data.

LUCIFER 0387 may be the same as 0388

acquire000125.fits failed to see the star. If the new algorithm finds these fainter stars reliably, then we should switch to that as default.

We saw "spiral in" a lot in doing these tests. Frequently it would fail because it exceeded the 6 cycles. These should probably keep trying if there is a star in the guide image.

Implement/Test New Pointing Model
Implemented, had large IE/CA offset because of the thermal issues?

02:34:17.635 Preset sent. 02:34:22 waiting for AZcam image. Killed and restarted gain

Focus LUCIFER/AGW on sky (30-45 minutes/field):
02:38:12.852 Preset sent again, different GS. luci_20090108_NNNN (saved as cubes of 10x2s)

393 0 Closed loop
394 -1500 Z4
395 -750 Z4
396 0 Z4
397 +750 Z4
398 +1500 Z4
399 0 Z4

400 0 Closed loop
401 +1500 Z4
402 +750 Z4
403 0 Z4
404 -750 Z4
405 -1500 Z4
406 0 Z4

AGW nominal focus offset by 3.75mm (600nm)

For -3.75mm from nominal

0 0463 -1500 0464 -750 0465 0 0466
  1. 0467 1500 0468 0 0469

We need to ADD 3.75mm to the nominal AGW focus.

Collected more collimation data for the remainder of the night.

FILE OBJECT TELALT M1-X M1-Y M1-Z M1RX M1RY M1RZ M1CTEMP
luci_20090108_0475.fits C10-12 38.156120 0.525118 -0.499834 1.056157 -40.688200 3.615658 0.000000 1.487000
luci_20090108_0476.fits C09-21 18.834599 0.605365 -0.055004 1.511773 -48.563060 12.242850 0.000000 1.429667
luci_20090108_0477.fits C10-01 45.335655 0.273113 -0.386956 0.801396 -37.284030 3.103519 0.000000 1.380333
luci_20090108_0478.fits C10+13 52.368435 -0.022868 -0.225811 0.576148 -34.838620 3.824807 0.000000 1.345333
luci_20090108_0479.fits C09+27 43.890667 0.373858 -0.419967 0.865003 -38.013410 2.311683 0.000000 1.296667
luci_20090108_0480.fits C10+26 57.323124 -0.165083 -0.087349 0.396570 -33.668730 3.620266 0.000000 1.256667
luci_20090108_0481.fits C13+10 66.985497 -0.638093 0.209029 0.008169 -32.942170 6.383794 0.000000 1.217000
luci_20090108_0482.fits C13+20 76.702698 -1.081612 0.649618 -0.394599 -33.263620 9.566458 0.000000 1.183333
luci_20090108_0483.fits C13+25 81.275200 -1.241312 0.815639 -0.586707 -34.817310 10.689030 0.000000 1.119667
luci_20090108_0484.fits C13+15 71.571579 -0.820572 0.322957 -0.198393 -33.981300 7.265231 0.000000 1.097667
luci_20090108_0485.fits C13+05 61.545578 -0.380532 -0.015763 0.184536 -33.937680 4.478889 0.000000 1.083333
luci_20090108_0486.fits C13-05 51.460659 0.028943 -0.269645 0.513355 -35.698280 2.836664 0.000000 1.065333

06:54 WFS no longer correcting, too bright. Closing.

-- DavidThompson - 07 Jan 2009
Topic revision: r18 - 18 Jan 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback